# Survey on Factuality in Large Language Models: Knowledge, Retrieval and Domain-Specificity

### v1

@@ 3.2 Benchmarks for Factuality Evaluation
HaluEval (Li et al., 2023c) is a benchmark designed to understand and evaluate the propensity of LLMs like ChatGPT to generate hallucinations. A hallucination, in this context, refers to content that either conflicts with the source or cannot be verified based on factual knowledge. The HaluEval benchmark offers a vast collection of generated and human-annotated hallucinated samples, aiming to evaluate the performance of LLMs in recognizing such hallucinations. The benchmark utilizes a two-step framework, termed ”sampling-then-filtering”, based on ChatGPT to generate these samples. Additionally, human labelers were employed to annotate hallucinations in ChatGPT responses. The HaluEval benchmark is a comprehensive tool that not only evaluates the hallucination tendencies of LLMs but also provides insights into the types of content and the extent to which these models are prone to hallucinate.
@@
Huang et al. (2023b) propose C-Eval, the first comprehensive Chinese evaluation suite. It can be used to evaluate the advanced knowledge and reasoning abilities of foundational models within a Chinese context. The evaluation suite comprises multiple-choice questions spanning 52 diverse disciplines, with four levels of difficulty: middle school, high school, college, and professional. Additionally, C-Eval Hard is introduced for very challenging subjects within the C-Eval suite, which demands advanced reasoning abilities to solve. Their evaluations of state-of-the-art LLMs, including both English and Chinese-oriented models, have shows that there is significant room for improvement as only GPT-4 managed to achieve an average accuracy of over 60%. The authors focus on assessing LLMs’ advanced abilities within a Chinese context. The researchers assert that LLMs intended for a Chinese environment should be evaluated based on their knowledge of Chinese users’ primary interests, such as Chinese culture, history, and laws. With C-Eval, the authors aim to guide developers in understanding the abilities of their models from multiple dimensions to facilitate the development and growth of foundational models for Chinese users. Simultaneously, C-Eval has not just introduced a whole suite but subsets that can serve as individual benchmarks, thereby assessing certain model abilities and analyzing key strengths and limitations of foundational models. Experiment results have shown that although GPT-4, ChatGPT, and Claude were not exclusively tailored for Chinese data, they emerged as the top performers on C-Eval.
@@
The Pinocchio benchmark (Hu et al., 2023) serves as an extensive evaluation platform, emphasizing factuality and reasoning for LLMs. This benchmark encompasses 20,000 varied factual queries from diverse sources, timeframes, fields, regions, and languages. It tests an LLM’s capability to discern combined facts, process both organized and scattered evidence, recognize the temporal evolution of facts, pinpoint minute factual disparities, and withstand adversarial inputs. Each reasoning challenge within the benchmark is calibrated for difficulty to allow for detailed analysis.
@@
Vu et al. (2023) present a dynamic benchmark called FreshQA designed to evaluate up-to-date world knowledge of LLMs. Its questions range from those that are never-changing to those that are fast-changing, as well as questions based on false premises. The aim is to challenge the static nature of LLMs and test their adaptability to ever-changing knowledge through human evaluations. They develop a reliable evaluation protocol that uses a two-mode system: RELAXED and STRICT for a comprehensive understanding of model performance, ensuring that answers are confident, definitive, and accurate. They also provide a strong baseline named FRESHPROMPT, which seeks to enhance LLM performance by integrating real-time data from search engines. Initial experiments reveal that, outdated training data weakens the performance of LLMs and the FRESHPROMPT method can significantly enhance it. The research underscores the need for LLMs to be refreshed with current information to ensure their relevance and accuracy in a constantly evolving world.
@@
Several benchmarks, such as BigBench (Srivastava et al., 2023) and C-Eval (Huang et al., 2023b), encompass subsets that extend beyond the realm of factual knowledge or factuality. In this work, we specifically emphasize and focus on those subsets related to factuality.
There are benchmarks primarily designed for Pre-trained Language Models (PLMs) that can also be adapted for LLMs. Some studies use them for evaluating LLM’s factuality, but they are not so widely-used, so we have chosen to exclude them from the table for clarity. These benchmarks predominantly encompass knowledge-intensive tasks, as highlighted by (Petroni et al., 2021). Those benchmarks include NaturalQuestions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQ) (Joshi et al., 2017), OTT-QA (Chen et al., 2021b), AmbigQA Min et al. (2020) and WebQuestion (WQ) (Berant et al., 2013) of Open-domain Question Answering (QA) task, the HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), IIRC (Ferguson et al., 2020), MuSiQue (Trivedi et al., 2022) of multi-step QA, the ELI5 (Fan et al., 2019) of the Long-form QA task, the FEVER (Thorne et al., 2018), FM2 (Eisenschlos et al., 2021), HOVER (Jiang et al., 2020), FEVEROUS (Aly et al., 2021) of the Fact Checking task, the T-REx (Elsahar et al., 2018), zsRE (Levy et al., 2017) and LAMA (Petroni et al., 2019a) to examine factual knowledge contained in pretrained language models, the WikiBio (Lebret et al., 2016) of biography generation task, the RoSE (Liu et al., 2023c), WikiAsp (Hayashi et al., 2021) of summarization task, the KILT (Petroni et al., 2021) of comprehensive knowledge intensive tasks, the MassiveText (Rae et al., 2022b), Curation Corpus (Curation, 2020), Wikitext103 (Merity et al., 2016), Lambada (Paperno et al., 2016), C4 (Raffel et al., 2020b), Pile (Gao et al., 2020) of langauge modeling task, the WoW (Dinan et al., 2019), DSTC7 track2 (Galley et al., 2019a), DSTC11 track5 (Zhao et al., 2023a) of dialogue task, the RealToxicityPrompts Gehman et al. (2020) of toxicity reduction, the CommaQA (Khot et al., 2022), StrategyQA (Geva et al., 2021a), TempQuestions (Jia et al., 2018), IN-FOTABS (Gupta et al., 2020a) of diverse reasoning tasks.
@@
3.3 Factuality Evaluation Studies
TABLE V: Factuality Evaluation of LLMs. This table lists the evaluation metrics used by studies aiming to improve the factuality of LLMs. The column ‘Human Eval’ denotes if the evaluation incorporated human evaluation. The ‘Granularity’ column specifies the level of evaluation, with token-level (T) and sentence-level (S) distinctions.
Reference Task Dataset Metrics Human Eval Evaluated LLMs Granularity
Retro
(Borgeaud et al., 2022)
QA,
Language
Modeling
MassiveText,
Curation Corpus,
Wikitext103,
Lambada,
C4,Pile, NQ
PPL,
ACC,
Exact Match
Retro
T
GenRead
(Yu et al., 2023)
QA,
Dialogue,
Fact Checking
NQ, TQ, WebQ,
FEVER,
FM2, WoW
EM, ACC,
F1, Rouge-L
GPT3.5, Codex
GPT-3, Gopher
FLAN, GLaM
PaLM
S
GopherCite
(Menick et al., 2022b)
Self-supported QA
NQ, ELI5,
TruthfulQA
(Health
Law, Fiction
Conspiracies)
Human Score
GopherCite
S
(Trivedi et al., 2023)
QA
HotpotQA, IIRC
2WikiMultihopQA,
MuSiQue(music)
Retrieval recall,
Answer F1
GPT-3
FLAN-T5
S/ T
(Peng et al., 2023)
QA,
Dialogue
DSTC7 track2,
DSTC11 track5,
OTT-QA
ROUGE, chrF,
BERTScore,
Usefulness,
Humanness…
ChatGPT
S/ T
CRITIC
(Gou et al., 2023)
QA
Toxicity Reduction
AmbigNQ,
TriviaQA,
HotpotQA,
RealToxicityPrompts
Exact Match,
maximum toxicity,
perplexity,
n-gram diversity,
AUROC…,
GPT-3.5
ChatGPT
T
(Khot et al., 2023)
QA,
long-context
QA
CommaQA-E,
2WikiMultihopQA,
MuSiQue,
HotpotQA
Exact Match,
Answer F1
GPT-3
FLAN-T5
T
ReAct
(Yao et al., 2023a)
QA
Fact Verification
HotpotQA,
FEVER
Exact Match,
ACC
PaLM
GPT-3
S/T
(Jiang et al., 2023b)
QA,
Commonsense
Reasoning,
long-form QA…
2WikiMultihopQA,
StrategyQA,
ASQA, WikiAsp
Exact Match,
Disambig-F1,
ROUGE,
entity F1…
GPT-3.5 T
(Lee et al., 2022b)
Open-ended
Generation
FEVER
Entity score,
Entailment
Ratio, ppl…
Megatron-LM T
SAIL
(Luo et al., 2023a)
QA
Fact Checking
UniLC
ACC
F1
LLaMA Vicuna
SAIL
T
(He et al., 2022)
Commonsense
Reasoning,
Temporal
Reasoning,
Tabular
Reasoning
StrategyQA,
TempQuestions,
IN-FOTABS
ACC GPT-3 T
(Pan et al., 2023a) Fact Checking
HOVER
FEVEROUS-S
Macro-F1
Codex
FLAN-T5
S
(Du et al., 2023)
Biography
MMLU
Unnamed
Biography
Dataset,
MMLU
ChatGPT
Evaluator,
ACC
Bard
ChatGPT
S
@@ 3.3 Factuality Evaluation Studies
Manakul et al. (2023) use an evaluation process that encompasses several key steps. Initially, synthetic Wikipedia articles are generated using GPT-3, focusing on individuals from the Wikibio dataset. Subsequently, manual annotation is performed at the sentence level, classifying sentences as “Major Inaccurate” , “Minor Inaccurate”, or “Accurate”, with “Major Inaccurate” denoting sentences unrelated to the topic. Passage-level scores are derived by averaging sentence-level labels, and identifying cases of total inaccuracies through score distribution analysis. Inter-annotator agreement is assessed using Cohen’s $\kappa$  Cohen (1960). Evaluation metrics primarily employ precision-recall curves (PR-Curves), distinguishing between “Non-Factual Sentences,” “Non-Factual* Sentences” (a specific subset), and “Factual Sentences.” These PR-Curves elucidate the trade-off between precision and recall for different detection methods.
Wang et al. (2023a) ask several LLMs, including ChatGPT, GPT-4 (OpenAI, 2023), BingChat (Microsoft, 2023) to answer open questions from NaturalQuestions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017). They manually estimate the accuracy of those LLMs on open question answering, and find that though LLMs can achieve nice performance but still far away from perfect.
@@
Chern et al. (2023) develop FACTOOL, which can be used as a factuality detector, to audit generative chatbots and evaluate the reliability of the responses they generate. Modern chatbots including GPT-4, ChatGPT, Claude, Bard (Google, 2023), and Vicuna (Chiang et al., 2023) are evaluated using FACTOOL, which is powered by GPT-4. To carry out the factuality evaluation on these chatbots, the team selected 30 prompts related to knowledge-based question answering (KB-QA), 10 code prompts, 10 math prompts, and 10 scientific prompts. The KB-QA prompts are collected from a previous study, code prompts from HumanEval, math prompts from yet another study, and the scientific prompts are generated by the authors themselves. Both claim-level and response-level accuracies are reported for each chatbot, with a weighted claim-level accuracy used to provide a more holistic and fair assessment of each chatbot’s factual accuracy. This weight is determined by the ratio of the number of prompts in each scenario. Interestingly, the results demonstrate that GPT-4 have the best weighted claim-level factual accuracy and response-level accuracy among all the evaluated chatbots. It is also observed that supervised fine-tuned chatbots, like Vicuna-13B, perform reasonably well in common scenarios like KB-QA but not as well in more challenging scenarios such as math, code, and scientific.
@@
In the realm of factuality evaluation, the Model Editing task holds a unique position, focusing on refining the internal knowledge of models. This task comes with its own set of specialized evaluation metrics. Predominantly, the Zero-Shot Relation Extraction (zsRE) Levy et al. (2017) and CounterFact Meng et al. (2022) serve as the primary benchmarks for assessing Model Editing techniques. When evaluating these methods, several key criteria emerge: Reliability: Post-editing, the model should consistently generate the intended output. Generalization: The model should be adept at producing the target output even when presented with paraphrased inputs. Locality: Edits should be localized, ensuring that facts not related to the specific edit remain intact. However, given the intricate web of interconnected facts, recent studies Zhong et al. (2023b); Yao et al. (2023b); Cohen et al. (2023a) have advocated for a more holistic evaluation approach. They introduce broader criteria for fact updates, encompassing aspects like portability, logical generalization, among others.
@@
Some work’s main contributions lie at the methods to improve the factuality of LLMs, and their evaluation part may also be informative and important when people reach the related study. So we choose to list evaluation parts in the Table V but not to discuss their evaluation in detail.
3.4 Evaluating Domain-specific Factuality

### v3

@@ 3.2. Benchmarks for Factuality Evaluation
HaluEval (Li et al., 2023a) is a benchmark designed to understand and evaluate the propensity of LLMs like ChatGPT to generate hallucinations. A hallucination, in this context, refers to content that either conflicts with the source or cannot be verified based on factual knowledge. The HaluEval benchmark offers a vast collection of generated and human-annotated hallucinated samples, aiming to evaluate the performance of LLMs in recognizing such hallucinations. The benchmark utilizes a two-step framework, termed ”sampling-then-filtering”, based on ChatGPT to generate these samples. Additionally, human labelers were employed to annotate hallucinations in ChatGPT responses. The HaluEval benchmark is a comprehensive tool that not only evaluates the hallucination tendencies of LLMs but also provides insights into the types of content and the extent to which these models are prone to hallucinate.
@@
Huang et al. (2023a) propose C-Eval, the first comprehensive Chinese evaluation suite. It can be used to evaluate the advanced knowledge and reasoning abilities of foundational models within a Chinese context. The evaluation suite comprises multiple-choice questions spanning 52 diverse disciplines, with four levels of difficulty: middle school, high school, college, and professional. Additionally, C-Eval Hard is introduced for very challenging subjects within the C-Eval suite, which demands advanced reasoning abilities to solve. Their evaluations of state-of-the-art LLMs, including both English and Chinese-oriented models, have shows that there is significant room for improvement as only GPT-4 managed to achieve an average accuracy of over 60%. The authors focus on assessing LLMs’ advanced abilities within a Chinese context. The researchers assert that LLMs intended for a Chinese environment should be evaluated based on their knowledge of Chinese users’ primary interests, such as Chinese culture, history, and laws. With C-Eval, the authors aim to guide developers in understanding the abilities of their models from multiple dimensions to facilitate the development and growth of foundational models for Chinese users. Simultaneously, C-Eval has not just introduced a whole suite but subsets that can serve as individual benchmarks, thereby assessing certain model abilities and analyzing key strengths and limitations of foundational models. Experiment results have shown that although GPT-4, ChatGPT, and Claude were not exclusively tailored for Chinese data, they emerged as the top performers on C-Eval.
@@
The Pinocchio benchmark (Hu et al., 2023a) serves as an extensive evaluation platform, emphasizing factuality and reasoning for LLMs. This benchmark encompasses 20,000 varied factual queries from diverse sources, timeframes, fields, regions, and languages. It tests an LLM’s capability to discern combined facts, process both organized and scattered evidence, recognize the temporal evolution of facts, pinpoint minute factual disparities, and withstand adversarial inputs. Each reasoning challenge within the benchmark is calibrated for difficulty to allow for detailed analysis.
@@
Several benchmarks, such as BigBench (Srivastava et al., 2023) and C-Eval (Huang et al., 2023a), encompass subsets that extend beyond the realm of factual knowledge or factuality. In this work, we specifically emphasize and focus on those subsets related to factuality.
There are benchmarks primarily designed for Pre-trained Language Models (PLMs) that can also be adapted for LLMs. Some studies use them for evaluating LLM’s factuality, but they are not so widely-used, so we have chosen to exclude them from the table for clarity. These benchmarks predominantly encompass knowledge-intensive tasks, as highlighted by (Petroni et al., 2021). Those benchmarks include NaturalQuestions (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQ) (Joshi et al., 2017), OTT-QA (Chen et al., 2021a), AmbigQA (Min et al., 2020) and WebQuestion (WQ) (Berant et al., 2013) of Open-domain Question Answering (QA) task, the HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), IIRC (Ferguson et al., 2020), MuSiQue (Trivedi et al., 2022) of multi-step QA, the ELI5 (Fan et al., 2019) of the Long-form QA task, the FEVER (Thorne et al., 2018), FM2 (Eisenschlos et al., 2021), HOVER (Jiang et al., 2020), FEVEROUS (Aly et al., 2021) of the Fact Checking task, the T-REx (Elsahar et al., 2018), zsRE (Levy et al., 2017) and LAMA (Petroni et al., 2019a) to examine factual knowledge contained in pretrained language models, the WikiBio (Lebret et al., 2016) of biography generation task, the RoSE (Liu et al., 2023a), WikiAsp (Hayashi et al., 2021) of summarization task, the KILT (Petroni et al., 2021) of comprehensive knowledge intensive tasks, the MassiveText (et al., [n. d.]), Curation Corpus (Curation, 2020), Wikitext103 (Merity et al., 2016), Lambada (Paperno et al., 2016), C4 (Raffel et al., 2020b), Pile (Gao et al., 2020) of langauge modeling task, the WoW (Dinan et al., 2019), DSTC7 track2 (Galley et al., 2019a), DSTC11 track5 (Zhao et al., 2023a) of dialogue task, the RealToxicityPrompts (Gehman et al., 2020) of toxicity reduction, the CommaQA (Khot et al., 2022), StrategyQA (Geva et al., 2021a), TempQuestions (Jia et al., 2018), IN-FOTABS (Gupta et al., 2020a) of diverse reasoning tasks.
@@
3.3. Factuality Evaluation Studies
Table 6. Factuality Evaluation of LLMs. This table lists the benchmarks that used in the enhancement work. The column ‘Human Eval’ denotes if the evaluation incorporated human evaluation. The ‘Granularity’ column specifies the level of evaluation, with token-level (T) and sentence-level (S) distinctions.
Reference Task Dataset Metrics Human Eval Evaluated LLMs Granularity Retro (Borgeaud et al., 2022) QA, Language Modeling MassiveText, Curation Corpus, Wikitext103, Lambada, C4,Pile, NQ PPL, ACC, Exact Match Retro T GenRead (Yu et al., 2023) QA, Dialogue, Fact Checking NQ, TQ, WebQ, FEVER, FM2, WoW EM, ACC, F1, Rouge-L GPT3.5, Codex GPT-3, Gopher FLAN, GLaM PaLM S GopherCite (Menick et al., 2022b) Self-supported QA NQ, ELI5, TruthfulQA (Health Law, Fiction Conspiracies) Human Score GopherCite S Trivedi et al. (2023) QA HotpotQA, IIRC 2WikiMultihopQA, MuSiQue(music) Retrieval recall, Answer F1 GPT-3 FLAN-T5 S/ T Peng et al. (2023) QA, Dialogue DSTC7 track2, DSTC11 track5, OTT-QA ROUGE, chrF, BERTScore, Usefulness, Humanness… ChatGPT S/ T CRITIC (Gou et al., 2023) QA, Toxicity Reduction AmbigNQ, TriviaQA, HotpotQA, RealToxicityPrompts Exact Match, maximum toxicity, perplexity, n-gram diversity, AUROC…, GPT-3.5 ChatGPT T Khot et al. (2023) QA, long-context QA CommaQA-E, 2WikiMultihopQA, MuSiQue, HotpotQA Exact Match, Answer F1 GPT-3 FLAN-T5 T ReAct (Yao et al., 2023b) QA, Fact Verification HotpotQA, FEVER Exact Match, ACC PaLM GPT-3 S/T Jiang et al. (2023a) QA, Commonsense Reasoning, long-form QA… 2WikiMultihopQA, StrategyQA, ASQA, WikiAsp Exact Match, Disambig-F1, ROUGE, entity F1… GPT-3.5 T Lee et al. (2022b) Open-ended Generation FEVER Entity score, Entailment Ratio, ppl… Megatron-LM T SAIL (Luo et al., 2023a) QA, Fact Checking UniLC ACC F1 LLaMA Vicuna SAIL T He et al. (2022) Commonsense Reasoning, Temporal Reasoning, Tabular Reasoning StrategyQA, TempQuestions, IN-FOTABS ACC GPT-3 T Pan et al. (2023c) Fact Checking HOVER FEVEROUS-S Macro-F1 Codex FLAN-T5 S Du et al. (2023) Biography, MMLU Unnamed Biography Dataset, MMLU ChatGPT Evaluator, ACC Bard ChatGPT S Asai et al. (2023) QA, CItation Generation PopQA, Pubhealth, ASQA, ARC, Unamed Biography ACC, Factscore, Exact Match, ROUGE, MAUVE… LLaMA2 ChatGPT Perplexity.ai Alpaca SAIL S/T
@@ 3.3. Factuality Evaluation Studies
Manakul et al. (2023) use an evaluation process that encompasses several key steps. Initially, synthetic Wikipedia articles are generated using GPT-3, focusing on individuals from the Wikibio dataset. Subsequently, manual annotation is performed at the sentence level, classifying sentences as “Major Inaccurate” , “Minor Inaccurate”, or “Accurate”, with “Major Inaccurate” denoting sentences unrelated to the topic. Passage-level scores are derived by averaging sentence-level labels, and identifying cases of total inaccuracies through score distribution analysis. Inter-annotator agreement is assessed using Cohen’s $\kappa$  (Cohen, 1960). Evaluation metrics primarily employ precision-recall curves (PR-Curves), distinguishing between “Non-Factual Sentences,” “Non-Factual* Sentences” (a specific subset), and “Factual Sentences.” These PR-Curves elucidate the trade-off between precision and recall for different detection methods.
BSChecker (Hu et al., 2023b) is a factual errors detection framework that detects inaccuracies on triplet-level granularity. Its pipeline comprises a claim extractor and a checker. The claim extractor extracts triplets from model’s response as claims formed with (subject, predicate, object). Meanwhile, the checker compare these claims triplets with references, categorizing each triplet to one of three hallucination labels: Entailment, Contradiction or Neutral, drawing inspiration from the Natural Language Inference task. BSChecker offers a benchmark that distinguishes three task settings: Zero Context, Noisy Context and Accurate Context covering tasks of closed-book QA, retrieval-augmented generation, summarization, closed-QA and information extraction.
@@
FACTOOL (Chern et al., 2023) is a tool designed to function as a factuality detector, with the primary purpose of auditing generative chatbots and assessing the reliability of their outputs. This tool is employed to evaluate several contemporary chatbots, including GPT-4, ChatGPT, Claude, Bard (Google, 2023), and Vicuna (Chiang et al., 2023). Notably, FACTOOL itself leverages the capabilities of GPT-4. For the evaluation process, the researchers have curated a diverse set of prompts: 30 from knowledge-based question answering (KB-QA), 10 each from code, math, and scientific domains. The KB-QA prompts were sourced from a prior study, code prompts were taken from HumanEval, math prompts from another distinct study, while the scientific prompts were crafted by the authors themselves. The evaluation metrics included both claim-level and response-level accuracies for each chatbot. To offer a more comprehensive and equitable evaluation, a weighted claim-level accuracy is used. The weighting is determined based on the proportion of prompts from each category. The findings are illuminating. GPT-4 emerge as the top performer in terms of both weighted claim-level factual accuracy and response-level accuracy among all the chatbots assessed. Another intriguing observation is that chatbots that underwent supervised fine-tuning, such as Vicuna-13B, exhibited commendable performance in standard scenarios like KB-QA. However, their performance dip in more intricate scenarios, including those involving math, code, and scientific queries.
Wang et al. (2023b) ask several LLMs, including ChatGPT, GPT-4 (OpenAI, 2023), BingChat (Microsoft, 2023) to answer open questions from NaturalQuestions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017). They manually estimate the accuracy of those LLMs on open question answering, and find that though LLMs can achieve nice performance but still far away from perfect. Besides, they evaluate whether the GPT-3.5 can assess the correctness of LLM-generated responses, and find negative results, even if the golden answer is also presented. Similarity, Fu et al. (2023a) ask LLMs, such as GPT-2 and GPT-4, to directly score the factuality of a summary, and find no significant correlation between LLM’s factuality indicators and human evaluations.
@@
CONNER (Chen et al., 2023b), a framework that evaluates LLMs as generators of knowledge. It focuses on six areas: Factuality, Relevance, Coherence, Informativeness, Helpfulness, and Validity. It evaluates whether the generated information can be backed by external proof (Factuality), is relevant to the user’s query (Relevance), and is logically consistent (Coherence). It also checks if the knowledge provided is novel or surprising (Informativeness). The Extrinsic evaluation measures whether the knowledge enhances downstream tasks (Helpfulness) and its results are factually accurate (Validity).
@@
In the realm of factuality evaluation, the Model Editing task holds a unique position, focusing on refining the internal knowledge of models. This task comes with its own set of specialized evaluation metrics. Predominantly, the Zero-Shot Relation Extraction (zsRE) (Levy et al., 2017) and CounterFact (Meng et al., 2022) serve as the primary benchmarks for assessing Model Editing techniques. When evaluating these methods, several key criteria emerge: Reliability: Post-editing, the model should consistently generate the intended output. Generalization: The model should be adept at producing the target output even when presented with paraphrased inputs. Locality: Edits should be localized, ensuring that facts not related to the specific edit remain intact. However, given the intricate web of interconnected facts, recent studies Zhong et al. (2023b); Yao et al. (2023a); Cohen et al. (2023a) have advocated for a more holistic evaluation approach. They introduce broader criteria for fact updates, encompassing aspects like portability, logical generalization, among others.
RAGAS (Es et al., 2023) proposes a framework for the reference-free evaluation of Retrieval-Augmented Generation (RAG) systems in Large Language Models (LLMs). This framework primarily assesses the RAG system’s ability to identify relevant and key context paragraphs, the LLM’s fidelity in utilizing these paragraphs, and the overall quality of the generated content. The framework focuses on three aspects of quality: Faithfulness: The generated answers should be based on the given context. To evaluate faithfulness, the generated content is broken down into sentences, which are then transformed into one or more short assertions. Each assertion’s faithfulness is assessed against the retrieved content, with the faithfulness score being the ratio of faithful assertions to the total number of assertions. Answer Relevance: The generated answers should appropriately address the posed questions. This is evaluated by having the LLM generate potential questions from each piece of generated content. An embedding model then calculates the similarity between each potential question and the original question. The answer relevance score is the average of these similarities. Context Relevance: The retrieved context should be highly relevant, containing as little irrelevant information as possible. This is assessed by using the LLM to extract sentences from the retrieved content that are relevant to the question. The context relevance score is then calculated based on the proportion of these relevant sentences to the total number of sentences retrieved. The authors created a WikiEval dataset, consisting of question-context-answer triplets with human judgments. This allows for the measurement of how closely different evaluation frameworks align with human assessments of faithfulness, answer relevance, and context relevance. The paper compares two baselines: GPTScore, which requires ChatGPT to rate the three quality dimensions on a scale of 0 to 10, and GPT Ranking, which does not require ChatGPT to choose a preferred answer/context but includes definitions of the considered quality metrics in the prompt. When evaluating answer relevance, a specific prompt is used. Overall, the RAGAS framework proposed in the paper aligns more closely with human judgment compared to the two baselines.
@@
Some work’s main contributions lie at the methods to improve the factuality of LLMs, and their evaluation part may also be informative and important when people reach the related study. So we choose to list evaluation parts in the Table 6 but not to discuss their evaluation in detail.
3.4. Evaluating Domain-specific Factuality